Skip to content

OSG Technology Area Meeting, 7 May 2018

Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Carl, Derek, Edgar, Marian, Suchandra, TimT

Announcements

BrianL and Derek out Thursday afternoon and Friday

Triage Duty

  • This week: TimT
  • Next week: Carl
  • 11 (-3) open tickets
  • 10 (-2) open LCMAPS VOMS tickets

JIRA

# of tickets Δ State
110 -5 Open
40 -4 In Progress
24 +10 Ready for Testing
11 +7 Ready for Release

OSG Software Team

  • Items from last week:
    • Derek: Ticket exchanges with Freshdesk
    • Suchandra: Reversing ticket comment order in Freshdesk
    • BrianL: Spoke with Marina about adding a banner to ticket
  • Freshdesk usage discussion: tentatively Thursday 5/17
  • JIRA migration: soon (?) but no date set
  • Office hours: Tuesdays 2-3pm CDT and Thursdays 10-11am CDT

Discussion

  • Derek will test ticket exchange with HCC RT and Freshdesk Tuesday morning
  • Suchandra: reverse comment order requires plugins that apply across the board and may break other things. Won't pursue it.
  • Action item (BrianL): Look into ticket types for operations/software Freshdesk groups

Support Update

  • LIGO (Edgar) - XRootD file transfer failed due to OSG proxies being signed with an expired VOMS certificate
  • UMD (Suchandra) - Not transitioning off of LCMAPS VOMS until the summer. Should be fine as a CMS T3.

OSG Release Team

  • OSG 3.4.12 - Mostly Ready for Release
    • GlideinWMS 3.2.22
      • proxy renewal script properly process with the default FQAN
    • Gratia probes
      • Fixed Slurm probe bugs found at UFL
      • HTCondor probe: projectname case insensitive comparison
      • PBS probe: Handle raw seconds as well as hh:mm:ss for durations
      • Drop GRAM and glexec probe
    • XRootD 4.8.3 - in testing
  • OSG 3.3.34 - Ready for Release
    • Eradicate grid.iu.edu references
    • xrootd-lcmaps 1.2.1-3: fixed crashes on Enterprise Linux 6 when request were made using HTTPS
    • frontier-squid: fixed startup problem under SELinux

Discussions

None this week

OSG Investigations Team

GOC Transition is going to dominate the Investigations Team's time over the next couple of weeks. I want to give general status of the transitions we are participating in. So, I will give a couple of states:

  1. Have not started
  2. In development - Could be provisioning hardware, adapting software, installing software…
  3. Ready for testing - In a state where it can be tested by the consumers
  4. Deployed
Service Status
OIM In Development
OASIS Deployed
CVMFS Deployed
Software Repo Deployed
OSG Display Deployed
StashCache Redirector Deployed
OSG Collector Ready for testing

Changes:

  • OASIS: Testing -> Deployed: Transitioned on 5/3
  • CVMFS: Testing -> Deployed: Transitioned on 5/3
  • OSG Collector needs discussion. CE's do not transition to the new collector without a restart.

Ongoing

Discussions

None this week